### v3

# LLMEval-3: A Large-Scale Longitudinal Study on Robust and Fair Evaluation of Large Language Models

Abstract
Existing evaluation of Large Language Models (LLMs) on static benchmarks is vulnerable to data contamination and leaderboard overfitting, critical issues that obscure true model capabilities. To address this, we introduce LLMEval-3, a framework for dynamic evaluation of LLMs. LLMEval-3 is built on a proprietary bank of 220k graduate-level questions, from which it dynamically samples unseen test sets for each evaluation run. Its automated pipeline ensures integrity via contamination-resistant data curation, a novel anti-cheating architecture, and a calibrated LLM-as-a-judge process achieving 90% agreement with human experts, complemented by a relative ranking system for fair comparison. An 20-month longitudinal study of nearly 50 leading models reveals a performance ceiling on knowledge memorization and exposes data contamination vulnerabilities undetectable by static benchmarks. The framework demonstrates exceptional robustness in ranking stability and consistency, providing strong empirical validation for the dynamic evaluation paradigm. LLMEval-3 offers a robust and credible methodology for assessing the true capabilities of LLMs beyond leaderboard scores, promoting the development of more trustworthy evaluation standards. Our github repository is available at https://github.com/llmeval/llmeval-3.
1 Introduction
The rapid advancement of Large Language Models (LLMs) has led to a proliferation of benchmarks designed to assess their capabilities (DBLP:journals/corr/abs-2307-03109; DBLP:conf/emnlp/LaskarABRKKJB0P24; DBLP:journals/corr/abs-2506-11094). However, these benchmarks predominantly rely on a static evaluation paradigm where models are tested on fixed, publicly available datasets (DBLP:journals/corr/abs-2503-04149). This approach is fundamentally vulnerable to data contamination and test set overfitting, contributing to a growing “evaluation crisis” where benchmark scores may no longer reliably reflect a model’s true, generalizable abilities (DBLP:journals/corr/abs-2412-03597; DBLP:conf/naacl/Deng0TGC24; DBLP:journals/corr/abs-2402-02823). For example, GPT-4 achieved exact match rates of 52% and 57% when guessing the masked options in MMLU (mmlu) test sets, far exceeding random chance (leakage:deng2024). Similarly, Qwen-1.8B has been shown to exactly replicate complete n-grams from both training and test splits of GSM8K (GSM8K) and MATH (MATH), including 25 exact 5-gram matches in the MATH test set, indicating potential undetected data leakage and memorization (benchmarkleakage:xu2024). Furthermore, static benchmark designs, dataset contamination, and biased evaluation protocols could create misleading perceptions of LLM capabilities, undermining the reliability of current performance assessments (DBLP:journals/corr/abs-2412-03597).
The crisis compels us to shift our focus from what capabilities to evaluate, such as knowledge and reasoning, to a more foundational question: how to evaluate in a manner that is robust, fair, and resistant to strategic manipulation. Based on our analysis of the current evaluation crisis, we identify three fundamental challenges that must be addressed to construct a trustworthy evaluation framework. These challenges correspond to the core stages of any assessment: the data, the protocol, and the ranking system.
Figure 1: The LLMEval-3 framework comprises three core stages. First, in Data Construction, diverse exam data are collected, filtered by experts, standardized into original questions in JSON format, and expanded to enhance formality and diversity. Second, Fair Testing involves sampling 1,000 unique questions for each model’s evaluation, transmitting data via a secure JWT stream, and collecting model responses. Third, Fair Ranking entails selecting evaluators using Cohen’s $\kappa$ coefficient under human-machine agreement, generating relative rankings based on evaluation scores (0–3), validating system robustness through ablation studies (incorporating varied reasoning styles, question resampling, and volume adjustments), and analyzing five key dimensions of model errors to verify ranking fairness. Column abbreviations: Eng. (Engineering), Econ. (Economics), Edu. (Education), Lit. (Literature), Mgmt. (Management), Sci. (Science), Hist. (History), Med. (Medicine), Mil. (Military).
Challenge 1: How can we ensure the integrity of the evaluation data? The cornerstone of any benchmark is its test set. If the data is public or predictable, it becomes susceptible to leakage into model training sets, leading to inflated scores that reflect memorization rather than true capability. Therefore, building a benchmark that is inherently resilient to data contamination is the primary challenge. To address this, we construct a large-scale, private question bank of over 220k graduate-level questions, which are held privately and augmented with structural variations to prevent memorization.
Challenge 2: How can we design an unpredictable assessment protocol? Even with private data, a static protocol with fixed test sets can be reverse-engineered or exploited over time. A truly robust evaluation requires a dynamic and unpredictable process that prevents any form of strategic gaming. To achieve this, we implement a dynamic evaluation protocol where models are served unseen, randomly sampled questions in each session through a secure, two-layer anti-cheating architecture, ensuring that each evaluation is unique and unpredictable.
Challenge 3: How can we establish a fair and stable ranking system under dynamic conditions? When the evaluation content is not fixed, traditional absolute scoring becomes unreliable for cross-model comparison. A fair ranking system must remain stable and consistent, even when models are tested on different, albeit equivalent, sets of questions. To solve this, we develop a novel relative ranking system powered by a highly-calibrated LLM-as-a-Judge framework. This system establishes rankings by comparing model performance relative to each other within each evaluation session, rather than relying on absolute scores, ensuring fair and robust rankings that demonstrate negligible variance across different data samples.
Collectively, these solutions constitute LLMEval-3, a comprehensive framework for dynamic LLM evaluation. We leverage this framework to conduct an extensive longitudinal study on our official platform, spanning from the second half of 2023 to the first half of 2025. Throughout this period, we have continuously evaluated nearly 50 leading proprietary and open-source models, typically assessing them within one week of their public release. To ensure the fairness and timeliness of our evaluation, we constantly update our private question bank. Each model undergoes at least three independent, randomly sampled evaluation runs to ensure the stability of our findings. To date, this rigorous, ongoing effort has accumulated over 150k evaluation data points.
Leveraging this large-scale evaluation data, we conduct systematic investigations across three key research questions corresponding to our core challenges. Our empirical analysis reveals several key findings that provide new insights into LLM capabilities and evaluation practices. We summarize several key findings: we find that (a) all models converge to a performance ceiling around 90% with persistent gaps in specialized domains like literature and medicine, (b) dynamic rankings diverge significantly from static benchmarks, and static benchmarks suffer from severe data contamination, (c) our framework demonstrates exceptional ranking stability with negligible variance under multi-round resampling and varying sample sizes. More insightful findings are presented in Section 3.
Overall, our contributions are threefold:
  1. 1.
    We construct LLMEval-3, a large-scale anti-cheating evaluation platform featuring a proprietary 220k question bank, secure dynamic sampling protocols, and robust anti-manipulation mechanisms for trustworthy LLM assessment.
  2. 2.
    We conduct an extensive 20-month longitudinal evaluation campaign across nearly 50 leading models, accumulating over 150k evaluation data points through continuous anti-cheating assessments and comparative analysis with static benchmarks.
  3. 3.
    We conduct extensive empirical analysis across three research questions corresponding to our core challenges, revealing seven key findings about model performance ceilings, ranking stability, and contamination vulnerabilities in current evaluation practices.
2 Design
In this section, we detail the design and implementation of LLMEval-3, which addresses the three fundamental challenges through a three-stage framework. Dataset Construction tackles data integrity by building a contamination-resistant private question bank. Evaluation Process addresses unpredictable assessment through dynamic sampling and anti-cheating mechanisms. Ranking System implements a calibrated ranking system for fair model comparison.
2.1 Dataset Construction
To build a contamination-resistant and high-quality question bank, we sourced postgraduate and undergraduate exam questions from Chinese universities, covering 13 primary and over 50 secondary academic disciplines. Figure 1 offers a comprehensive overview of the design, highlighting three key stages: Dataset Construction, Fair Testing and Fair Ranking.
The construction follows a rigorous pipeline. First, we collect original exam questions from diverse formats and invite over 30 expert annotators for quality screening, which yields 78,009 high-quality original questions after eliminating those with factual errors or irrelevant answers. Second, we employ an LLM-driven augmentation process to expand coverage and diversity. For instance, each Multiple-Choice question with $n$ options is converted into $n$ Fill-in-the-Blank variants, while Material Analysis questions are decomposed into multiple true/false questions based on key information. Finally, all augmented questions undergo format verification and metadata enrichment to ensure quality and traceability.
As of early 2025, this process has resulted in the LLMEval-3 dataset, which comprises over 220k questions across six main categories. To maintain evaluation freshness and prevent contamination, we continuously expand the question bank through the same manual collection and automated augmentation pipeline. Detailed statistical analysis of the dataset, including disciplinary breakdown and content diversity, is provided in Appendix A.
2.2 Evaluation Process
To ensure a reliable and fair evaluation, we design a dynamic process centered on a multi-layered anti-cheating architecture. This approach guarantees that each evaluation is unique, robust against manipulation, and accurately reflects a model’s capabilities. The process is built upon two core strategies: dynamic question sampling and a secure delivery architecture.
Dynamic Question Sampling
To ensure unpredictability, each model evaluation is based on a unique set of 1,000 questions randomly sampled from our private question bank. A dynamic algorithm generates a unique, non-repeatable question sequence for each evaluation session, making it impossible to predict upcoming questions based on historical patterns. Furthermore, models are required to answer questions in the pre-allocated order, preventing any “cherry-picking” strategies. This dynamic sampling ensures that every evaluation is a distinct and isolated event, reflecting the model’s true generalization ability rather than its performance on a known set of problems.
Secure Anti-Cheating Architecture
The evaluation process is protected by a secure, two-layer anti-cheating architecture.
  • The Outer Layer (Access Control): This layer manages authentication and authorization. We use JSON Web Tokens (JWT)(rfc7519) to secure every API request, ensuring that only authenticated models can participate in an evaluation session. A strict Role-Based Access Control (RBAC) system prevents any cross-session or cross-user data access, isolating each evaluation.
  • The Inner Layer (Process Control): This layer enforces the evaluation rules. A multi-level quota system tracks the number of questions allocated, pending, and completed, effectively preventing models from attempting to acquire more questions than permitted or resubmitting answers. As a final safeguard, our system automatically strips all answers and explanations from the data transmitted to the model, ensuring that only the question content is exposed and preventing answer leakage through data parsing.
2.3 Ranking System
To establish fair and stable rankings under dynamic evaluation conditions, we develop a calibrated ranking framework that combines LLM-as-a-Judge evaluation with relative scoring mechanisms. This approach ensures consistent and reliable model comparisons even when different question sets are used across evaluation sessions.
LLM-as-a-Judge Evaluation
To quantify answer quality, we establish a standardized scoring metric with an integer range of [0,3], ensuring consistent evaluation of response efficacy across diverse model architectures. For scoring implementation, we uniformly employ GPT-4o (4o) as our judge, which has demonstrated high human-machine agreement through rigorous validation (detailed in Section 3.4). This choice of a single, validated scoring model eliminates potential biases introduced by varying evaluative criteria.
The scoring focuses on both core correctness and explanation quality, with core correctness serving as the primary indicator for score determination. The specific evaluation prompt and criteria are provided in Appendix B.
Evaluation Metrics
To mitigate systematic bias introduced by random sampling questions, LLMEval-3 employs both relative score and absolute scores as evaluation metrics.
The absolute score $S_{\text{model}}$ represents a model’s performance on $N=1000$ questions, where each question receives a score $s_{i}$ (with maximum score $s_{\text{max}}=3$), mapped to the [0,100] interval:
$S_{\text{model}}=\frac{\sum_{i=1}^{N}s_{i}}{N\times s_{\text{max}}}\times 100$ (1)
The relative score $R_{\text{SOTA}}^{\text{model}}$ is defined as the model’s absolute score relative to the current SOTA model’s absolute score on the same question set, mapped to the [0,100] interval:
$R_{\text{SOTA}}^{\text{model}}=\frac{S_{\text{model}}}{S_{\text{SOTA}}}\times 100$ (2)
In our current evaluation, we use Doubao-1.5-Thinking-Pro as the reference SOTA model for relative score calculations.
Model $R_{\text{SOTA}}^{\text{model}}$ $S_{\text{model}}$ Eng. Econ. Edu. Law Lit. Mgmt. Sci. Hist. Med. Mil.
Open-source LLMs
DeepSeek-R1 97.40 91.23 9.47 9.43 9.27 9.37 8.83 9.37 9.03 9.53 8.50 8.43
DeepSeek-V3 96.47 90.36 9.30 9.57 8.93 9.23 8.60 9.13 8.97 9.47 8.83 8.33
Qwen-3-235B 96.43 90.32 9.23 9.43 9.03 9.50 8.23 9.43 8.97 9.17 8.73 8.60
Qwen-3-32B 92.22 86.38 8.43 9.10 8.57 9.10 7.77 9.47 8.67 9.30 7.70 8.27
Closed-source LLMs
Doubao-1.5-Thinking-Pro 100.00 93.67 9.47 9.67 9.43 9.77 8.93 9.53 9.23 9.70 8.97 8.97
Gemini-2.5-Pro 97.22 91.07 9.20 9.47 9.20 9.30 8.43 9.63 9.07 9.40 8.50 8.87
Gemini-2.5-Pro-Thinking 97.15 91.00 9.13 9.50 9.37 9.47 8.40 9.63 9.20 9.27 8.30 8.73
Doubao-1.5-Pro 95.68 89.62 8.83 9.03 9.13 9.43 8.57 9.27 8.83 9.10 8.60 8.83
o1 93.36 87.45 8.90 9.30 8.67 8.77 7.73 9.27 8.90 8.97 8.17 8.77
Claude-Sonnet-4-Thinking 91.03 85.27 8.57 9.00 8.63 8.73 7.57 9.10 8.93 8.70 7.97 8.07
Claude-Sonnet-4 91.00 85.24 8.57 8.80 8.50 8.70 7.80 9.03 8.80 8.80 8.17 8.07
GPT-4o-search 89.40 83.74 8.27 8.77 8.43 8.67 7.77 8.80 8.20 8.73 8.27 7.83
GPT-4o 88.09 82.51 7.90 8.67 8.30 8.33 7.17 8.97 8.57 8.67 7.63 8.30
o3-mini 84.13 78.80 7.97 8.60 8.30 8.20 6.73 8.57 8.53 7.17 7.03 7.70
Table 1: Overall and Subject‐Level Scores. $R_{\text{SOTA}}^{\text{model}}$ represents the relative score (0-100 scale) as defined in Equation (2), with Doubao-1.5-Thinking-Pro as the reference SOTA model. $S_{\text{model}}$ represents the absolute score (0-100 scale) as defined in Equation (1). Subject-level scores use a 10-point scale.
3 Experiment and Analysis
Based on the three core challenges identified in our introduction, we design the LLMEval-3 evaluation framework. In this section, we systematically investigate three critical research questions that arise from these challenges:
Research Question I: What authentic capability distributions and longitudinal trends do LLMs exhibit under LLMEval-3?
Research Question II: How does LLMEval-3 dynamic evaluation compare with static benchmarks regarding ranking accuracy and contamination issues?
Research Question III: How stable and reliable is LLMEval-3’s relative ranking system under multi-round resampling and human-machine consistency validation?
Through comprehensive experiments designed to address these research questions, we aim to validate the effectiveness of our dynamic evaluation paradigm and provide empirical evidence for the superiority of contamination-resistant assessment frameworks.
3.1 Experimental Setup
Benchmarking LLMs on LLMEval-3
We tracked over 50 LLMs from late 2023 to mid-2025, presenting full results in Appendix C and focusing here on 13 representative models (proprietary and open-source). Each model is evaluated across three prompting paradigms (Zero-Shot, Few-Shot, Chain-of-Thought) and 10 academic disciplines.
Case Study and Error Analysis
We sample incorrect responses from all evaluated models and manually classified failures into five categories: disciplinary knowledge, misunderstanding, logical reasoning, factual inaccuracies, and format compliance.
Ablation Studies
We conduct comprehensive ablation studies across two key dimensions:
Benchmark Comparison: We measure Spearman correlation between LLMEval-3 and static benchmarks (AGIEval (AGIEval), C-Eval (c-eval)) and perform fill-in-the-blank replay tests (1,000 questions, three attempts each) to assess contamination.
Ranking Validation: We conduct multi-round resampling (n=1000, 2000, 4000) to test stability. Human-machine agreement validation involves two independent rounds of human evaluation across 13 models, with Cohen’s $\kappa$ coefficients computed against three LLM-as-Judge evaluators. Second, we run an ablation study comparing our relative ranking to the traditional Elo scoring system.
3.2 Research Question I
Finding 1: All models converge to a performance ceiling of around 90% over a longitudinal period, with leading open-source LLMs rivaling proprietary SOTA.
Figure 2 illustrates the performance growth trajectories of different model series over time. Table 1 presents comprehensive evaluation scores and subject-level breakdowns for each model under the LLMEval-3 framework. We can find that nearly all models approach a performance ceiling around 90-93% on academic knowledge tasks. This convergence indicates fundamental limits in current model architectures for knowledge-intensive evaluation.
Model ZS FS CoT Mean Variance
Claude-4-Sonnet-Thinking 85.27 85.60 86.87 85.91 0.48
DeepSeek-R1 91.23 89.33 88.43 89.67 1.36
Doubao-1.5-Thinking-Pro 93.67 90.63 91.73 92.01 1.57
Table 2: Capability under three prompting paradigms.
Figure 2: Trend of model series. Models of the same series are primarily illustrated in the same color for better distinction. The fitted curve highlights the overall growth of performance over the observed period.
Besides, the evaluation results show that top-performing models achieve remarkable scores: the proprietary Doubao-1.5-Thinking-Pro reaches 93.67 while the open-source DeepSeek-R1 achieves 91.23. These leading models substantially outperform established systems like GPT-4o at 82.51 and o3-mini at 78.80. This demonstrates that with sufficient scale and targeted fine-tuning, both proprietary and open-source LLMs can achieve comparable performance.
Finding 2: Models demonstrate significant domain-specific performance variations, with specialized ’thinking’ abilities offering only marginal gains.
Our analysis reveals a consistent pattern of domain-specific performance across all models. As shown in Table 1, models excel in technical and business domains like Management and Economics (achieving scores $\geq 9.40$) but consistently underperform in humanities and specialized fields such as Literature, Medicine, and Military (all scores below 9.0). This indicates robust mastery in some areas but also highlights persistent knowledge gaps. Further analysis into model variants shows that enabling specialized “thinking” modes provides only modest advantages. For example, Gemini-2.5-Pro-Thinking scores 91.00 overall (vs. 91.07 for the preview), while Claude-Sonnet-4-thinking outperforms its base version by just $\approx 0.03$ overall. These results suggest that while dedicated ’thinking’ optimizations offer slight cross-domain benefits, the primary performance differentiators remain rooted in the breadth of domain-specific knowledge, not enhanced reasoning modes.
Finding 3: In dynamic knowledge-intensive evaluations, prompting paradigms have minimal impact, whereas external augmentation significantly boosts performance.
As shown in Table 2, the performance variance across Zero-Shot (ZS), Few-Shot (FS), and Chain-of-Thought (CoT) prompts remains below 1.6 points for all evaluated models, indicating that core capabilities are not significantly influenced by the prompting format. Detailed evaluation results for the other models are provided in Appendix C. In stark contrast, enabling web search boosts GPT-4o’s overall score by +1.23 (from 82.51 to 83.74), with the largest gains in knowledge-heavy domains like Medicine (+0.64) and Literature (+0.60). This demonstrates that while models possess a stable internal knowledge base, augmenting it with external information is a highly effective strategy for enhancing performance.
Finding 4: Systematic error analysis reveals that disciplinary knowledge gaps and comprehension failures are the primary limitations of current models.
Our comprehensive analysis of incorrect responses identifies disciplinary knowledge errors (47.7%) and misunderstanding errors (39.8%) as the two most prevalent failure modes, accounting for nearly 90% of all incorrect responses. This distribution suggests that current limitations stem primarily from knowledge gaps and comprehension deficiencies rather than systematic reasoning failures. For instance, Figure 3 illustrates characteristic failures where models either exhibit a clear lack of domain-specific knowledge or fail to capture critical contextual cues like the word “first”, instead matching only superficial details. These cases demonstrate that while models possess sophisticated language understanding, they still struggle with deep domain expertise and nuanced contextual interpretation, which defines their current performance boundaries.
Figure 3: Distribution of model error causes and illustrative cases of the two most prevalent error types.
3.3 Research Question II
Finding 5: Dynamic rankings diverge significantly from static benchmarks, revealing a different landscape of model capabilities.
To quantify the difference between dynamic and static evaluation, we compare the model rankings from LLMEval-3 with those from two prominent static benchmarks, AGIEval and C-Eval. As shown in Table 3, all correlations are moderate ($\rho\approx 0.66\text{--}0.71$) and statistically insignificant ($p>0.1$), confirming that static evaluations yield markedly different model rankings compared to our dynamic, contamination-resistant framework.
Benchmark $\rho$ $p$-value
AGIEval (EN) 0.714 0.111
AGIEval (ZH) 0.657 0.156
C-Eval 0.657 0.156
Table 3: Spearman’s rank correlation between LLMEval-3 and static benchmarks.
Finding 6: Static Benchmarks Suffer from Severe Data Contamination.
To assess the difference in leakage risk between publicly available static benchmarks and our private question bank, we conduct fill-in-the-blank replay tests on AGIEval (EN), AGIEval (ZH), C-Eval. For each dataset, we attempt 1000 questions with three completion attempts each, counting a question as successfully recalled if at least two attempts are correct. Table 4 reports the number of successful completions for each model.
We observe that on public static benchmarks all models achieve substantially higher completion counts (up to 248), indicating extensive leakage of these questions into the training corpora. In contrast, on our private dataset the number of successful completions drops markedly (maximum 179), with most models near or below 100, demonstrating minimal memorization. This confirms that public static benchmarks suffer from serious data leakage risks, whereas our private question bank—being excluded from large-scale pretraining data—effectively mitigates such risks.
Model AGI (EN) AGI (ZH) C-Eval Ours
DeepSeek-V3 97 153 136 80
ClaudeSonnet-4 179 248 224 179
Doubao-1.5 66 105 117 76
o3-mini 58 74 45 75
GPT-4o 54 86 62 48
Qwen3-32B 55 77 72 36
Table 4: Comparison of successful fill-in completions for different models on static benchmarks and LLMEval-3.
3.4 Research Question III
Finding 7: The relative ranking system demonstrates exceptional stability, with negligible variance across multi-round resampling and varying sample sizes.
To validate the stability of the LLMEval-3 ranking system, we conduct rigorous stress tests involving multi-round resampling and varying sample sizes. In our resampling test, we draw two independent question samples and find that the model ranking order remains identical across both runs, with relative scoring exhibiting remarkably low variance (e.g., $\sigma^{2}=1.68$ for DeepSeek-V3 and just 0.25 for o3-mini). Furthermore, we test the system’s sensitivity to question volume by running evaluations with sample sizes of n=1000, 2000, and 4000. Across all volumes, the ranking order holds firm. These results, detailed in Appendix D, provide strong empirical evidence that our relative ranking system is exceptionally robust and reliable.
Figure 4: Cohen’s $\kappa$ coefficients measuring agreement between human evaluators and three LLM judges across evaluations. GPT-4o achieves almost perfect agreement with human judgments ($\kappa=0.907$).
Finding 8: The relative ranking system via LLM-as-Judge achieves high human-machine agreement and demonstrates superior robustness over alternative ranking systems.
To validate our relative ranking system, we assess it on two critical dimensions: human alignment and methodological robustness. First, we measure its agreement with human experts by conducting two rounds of human evaluation on 13 models and calculating Cohen’s $\kappa$ coefficient against our LLM-as-a-Judge. As shown in Figure 4, the system achieve near-perfect agreement with human judgments, confirming its reliability. Second, we conduct an ablation study comparing our relative ranking against a traditional absolute scoring system (Elo). The results in Figure 5 show that our relative ranking achieves higher Spearman’s $\rho$ correlation and greater stability across varying sample sizes. This dual validation provides strong evidence that our relative ranking system is both aligned with human evaluation and more robust than traditional ranking approaches.
4 Related Work
Static Knowledge-Intensive Benchmarks such as MMLU (mmlu) and C-Eval (c-eval), provided foundational insights by assessing factual knowledge, reasoning, and cross-domain generalization in language models. These benchmarks established important baselines for measuring model performance across diverse academic domains. However, their static question sets significantly amplify the risk of data leakage, causing inflated evaluation scores and misleading perceptions of model capabilities (DBLP:journals/corr/abs-2412-03597; leakage:xu2024). Moreover, repeated use of fixed datasets induces benchmark-specific overfitting, with models memorizing test answers rather than achieving genuine generalization (leakage:deng2024). LLMEval-1 and LLMEval-2 (llmeval:zhang2024) also fall into this static category. These fundamental vulnerabilities in static evaluation have motivated researchers to explore more dynamic, robust evaluation approaches that better reflect true model generalization abilities.
Dynamic Human-Preference Evaluation such as LMSYS (lmsys:zheng2024), Chatbot Arena (chatbotarena:chiang2024), and AlpacaEval (AlpacaEval) introduce dynamic evaluation settings by leveraging human preferences to compare model outputs in pairwise or multi-turn interactions. These approaches effectively capture conversational quality, coherence, and real-time performance, offering valuable insights beyond traditional static benchmarks. However, they often lack depth in evaluating domain-specific or complex reasoning tasks, and their reliance on non-expert crowdworkers can introduce bias and reduce reliability in assessments of specialized knowledge (humanprefer:raju2024). Furthermore, the focus on conversational abilities may not adequately assess systematic knowledge coverage across academic disciplines.
LLM-as-a-Judge has emerged as a promising paradigm for scalable evaluation, where language models act as evaluators of other models’ outputs. Notable works include G-Eval, which uses GPT-4 with chain-of-thought reasoning and form-filling paradigms to assess natural language generation quality (geval:liu2023), and MT-Bench-101, which evaluates multi-turn dialogues through pairwise comparisons with GPT-4 justifications (mtbench101). These approaches offer significant advantages including high scalability, rapid deployment, cost-effectiveness, and demonstrated alignment with human preferences. However, most existing efforts focus primarily on eliciting better scoring mechanisms, while systematic validation and calibration of the judges, along with mitigation of inherent biases, remain underexplored (judging:zheng2023). Our work addresses these gaps by providing comprehensive analysis and systematic validation of the LLM-as-a-judge paradigm with rigorous human-machine agreement studies.
Figure 5: Relative ranking consistently outperforms Elo, reaching near-perfect correlation as the subset grows.
5 Conclusion
We introduced LLMEval-3, a dynamic evaluation framework built on a private 220k-question bank, contamination-resistant pipeline, two-layer anti-cheating architecture, and LLM-as-Judge relative ranking to deliver fair, robust, and reproducible LLM assessments. Over a 20-month longitudinal study of nearly 50 open-source and proprietary models, we observed a consistent performance ceiling near 90%, persistent domain-specific gaps in literature, medicine, and military knowledge, and widespread data leakage in static benchmarks. Our relative ranking method showed negligible variance under multi-round resampling and varying sample sizes, while achieving near-perfect agreement with human experts. We further demonstrated that prompting format has minimal impact, whereas external web-search augmentation yields significant gains in knowledge-intensive tasks. These findings confirm the superiority of dynamic, contamination-resistant evaluation over static paradigms and highlight the need for more trustworthy benchmarking practices to guide future LLM development.
Appendix A Dataset
This section provides supplementary information on our LLMEval-3 dataset employed in the study, to clarify the academic disciplines and question types covered, and to elaborate methodologies for scaling the questions.
A.1 Categories of Academic Displines
Categories of Academic Disciplines
As illustrated in Figure 6, we collected graduate-level examination questions spanning 13 primary (Philosophical Sciences, Economic Sciences, Law, Education, Literature, History, Engineering, Agronomy, Medicine, Military Science, Management Sciences, Arts, and Sciences) and more than 50 secondary academic disciplines recognized by China’s Ministry of Education. Two-thirds of the questions are derived from Chinese universities’ Postgraduate Entrance Exams, and one-third are from Undergraduate Final Exams of comparable difficulty. Detailed distribution of question sources is listed in Table 5.
Categories of Question Types
The raw dataset encompasses a diverse range of original question formats, including multiple-choices, fill-in-the-blank, true-or-false, short-answer, term explanation, and material analysis questions. Following question expansion and formatting, we have unified all questions and their answers into a fill-in-the-blank-like question-answer format, abandoning the complex answer structures of various question types, such as options A to E in multiple-choice questions, and “true” or “false” in true-or-false questions. This natural question-answer format enables the dataset to more thoroughly showcase the model’s capabilities.
Type Number of Topics Proportion (%)
Undergraduate Final Exams 26633 34.1
Postgraduate Entrance Exams 51376 65.9
Total 78009 100.0
Undergraduate Final Exams 71038 31.1
Postgraduate Entrance Exams 157566 68.9
Total 228604 100.0
Table 5: Distribution of question number and proportions for Undergraduate Final Exams and Postgraduate Entrance Exams.
Figure 6: Categories of Primary and Secondary Academic Disciplines.
A.2 Details of Expanding the Dataset
As shown in Table 6, we have amassed a substantially large quantity of original questions, with a marked surge in numbers following the expansion of our dataset.
Subject Original Count Rewrited Count Increase Count Increase Percentage (%)
Philosophical Sciences 2194 10969 8775 399.95
Medicine 30772 109974 79202 257.38
Law 9262 30116 20854 225.16
Management Sciences 2448 7945 5497 224.55
Engineering 4926 13263 8337 169.24
Sciences 6182 15669 9487 153.46
Economic Sciences 9245 18124 8879 96.04
Military Science 611 1187 576 94.27
Education 2781 5094 2313 83.17
History 1749 3178 1429 81.70
Literature 7839 13085 5246 66.92
Total 78009 228604 150595 193.05
Table 6: Distribution of Original and Rewritten Counts Across Disciplines.
Original Data Construction Pipeline
Original questions, structured simply and organized by subject for initial collation, follow this construction pipeline: first, converting Excel, Word, and PDF test papers to TXT; then batch splitting into JSON-formatted questions via scripts; and finally conducting data screening.
The latter involves three steps: (1) Expert review removes factually erroneous or irrelevant questions. (2) Batch splitting classifies and isolates questions, addressing errors such as content overlap, missing questions, or misclassifications. (3) Format cleaning resolves encoding conflicts, special characters, symbol consistency, redundancy, and typos.
Data Expanding Pipeline
To accommodate diverse application scenarios, this study proposes an augmented data format that complements the original question structure, The augmented dataset incorporates comprehensive metadata, including primary disciplinary categories, secondary disciplinary categories, question descriptions, answer content, and unique identifiers (UUIDs), which facilitates categorized data management. An example of expanding the Multiple-Choices question is shown in Figure 7.
Two additional key verification procedures are implemented upon the core augmentation strategy: (1) Format validation, which entails checking the consistency of option counts for multiple-choice questions and the alignment of answer spaces for Fill-in-the-Blank questions; and (2) Redundancy checks, which involve detecting duplicates among split questions and ensuring the uniqueness of question UUIDs. The data formats before and after expanding is illustrated in Figure 8.
Original Multiple-Choices questions: Title: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most ( ). A. 39 B. 52 C. 111 D. 119 Answer: C Expanded questions: Question 1: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “39” in the provided space? Answer: False Question 2: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “52” in the provided space? Answer: False Question 3: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “111” in the provided space? Answer: True Question 4: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “119” in the provided space? Answer: False
Figure 7: Example of how to expand a Multiple Choice question.
category: Engineering, sub_category: Computer Science, question_uuid: a4b92619-f0ac-xxxx-xxxx-9f11ce3b4704, dialog: role: user, content: Judge the correctness of the following statements, answer true/false, and give your reasons. The continue statement in a loop breaks out of the current loop. hint: false, gpt4res: Answer: false. The reason is that the continue statement does not break out of the entire loop.It skips the rest of the current loop and immediately moves on to the next iteration of the loop instead of ending the current loop completely. gpt4judge: judge: Overall Rating: 3 The reason why you gave this rating: The LLM provided the correct answer (false) and gave a reasonable explanation. It accurately explained that the ‘continue‘ statement does not terminate the entire loop but instead skips the remaining part of the current iteration and proceeds to the next iteration of the loop. This explanation aligns with the correct behavior of the ‘continue‘ statement in programming., stars: 3, reason: The LLM provided the correct answer (false) and gave a reasonable explanation. It accurately explained that the ‘continue‘ statement does not terminate the entire loop but instead skips the remaining part of the current iteration and proceeds to the next iteration of the loop. This explanation aligns with the correct behavior of the ‘continue‘ statement in programming.
Figure 8: Detailed entries of a single question after expanding.
Appendix B Prompts
This section presents the complete set of prompts used in LLMEval-3 for different evaluation paradigms. We provide the specific prompt templates for few-shot learning, chain-of-thought reasoning, and LLM-based automated evaluation to ensure reproducibility of our experimental results.
The prompts used for testing few-shot and chain-of-thought methods are shown in Figures 9 and 10, respectively. The prompt used for LLM-based evaluation is shown in Figure 11.
Input: Here are several examples: Question: Please determine the correctness of the following statement. Answer with true/false and provide a reason. If all elements below the diagonal in the adjacency matrix of a directed graph are zero, then the graph must have a topological ordering. Answer: true. Because if all elements below the diagonal in the adjacency matrix of a directed graph are zero, the graph is a Directed Acyclic Graph (DAG), so it must have a topological ordering. Question: Please explain the following term: “Double Hundred Policy”. Answer: It refers to “let a hundred flowers bloom, let a hundred schools of thought contend.” This was a policy on science and culture officially proposed by Mao Zedong in 1956 and confirmed by the Central Committee of the Communist Party of China. The policy was severely undermined after 1957, but was reestablished and implemented following the Third Plenary Session of the 11th Central Committee. Question: Please determine the correctness of the following statement. Answer with true/false and provide a reason. According to the convertibility theory, the scope of commercial banks’ assets expanded from short-term turnover loans to consumer loans. Answer: false. Convertibility Theory: also known as the asset conversion theory. This theory suggests that to maintain liquidity for withdrawals, commercial banks can invest part of their funds in transferable securities. Since these profitable assets can be sold at any time and converted into cash, loans are not necessarily limited to short-term and self-liquidating types. Clearly, this theory emerged in the context of developing financial instruments and markets. Significance: it expanded the scope of bank asset operations. Drawbacks: it does not guarantee that assets can be liquidated without capital loss (which requires high asset quality and stable market conditions); it is also constrained by central bank monetary policy (e.g., the risk of a rise in discount rates). This theory provides a theoretical basis for Chinese commercial banks to engage in securities business (investment operations). However, in China, commercial banks’ investment activities are restricted due to: 1. limited investment instruments and underdeveloped credit mechanisms; 2. management systems narrowing investment scopes, and separation of operations preventing commercial banks from investing in stocks; 3. the nature of state-owned commercial banks limits their willingness for autonomous investment. Question: Why can the results of animal experiments not be fully applied to clinical practice? Answer: Because there are differences between humans and animals not only in cellular morphology and metabolism, but also fundamentally due to the highly developed human nervous system, which is associated with language and thought (the second signaling system). Although there are similarities, the essential differences mean that human diseases cannot all be replicated in animals. Even if they can be replicated, animal responses are simpler than human responses. Therefore, results from animal experiments cannot be mechanically and fully applied to clinical practice without analysis. Only by comparing, analyzing, and synthesizing animal experiment results with clinical data can they be used as references in clinical medicine and provide a basis for studying the causes, mechanisms, prevention, and treatment of clinical diseases. The following is the question to be answered: Question:{question} Answer:
Figure 9: The Few-shot Prompt Template.
Input: {question} Please think step by step and provide the final answer.
Figure 10: The Chain-of-Thought Prompt Template.
Input: Please evaluate the following response from the LLM regarding a discipline-specific question based on the following criteria. You must score it on a scale of 0, 1, 2 or 3 stars: Overall Rating: 0 star indicates wrong answer with a wrong explanation 1 stars indicate wrong answer but a partially reasonable explanation 2 stars indicate a correct answer with a partially reasonable explanation 3 stars indicate an correct answer with a reasonable explanation User: {question} LLM: {LLM response} The correct answer to user’s question is: {correct answer} You must provide your feedback in the following format: “Overall Rating”:numbers of its stars(int) The reason why you gave this rating: ¡Your Reason¿(str)’
Figure 11: The Prompt Template for LLM Judgement.
Appendix C LLMEval-3 Leaderboard
This section presents comprehensive evaluation results from our longitudinal study tracking over 50 LLMs from late 2023 to mid-2025. We provide complete performance rankings and analyze the consistency of model capabilities across different prompting paradigms.
We tracked over 50 LLMs from late 2023 to mid-2025. Here, we present the complete evaluation results of our model assessments. The comprehensive results, including scores for all models across 10 academic disciplines, are presented in Table 7.
The models we selected in main paper was evaluated across three prompting paradigms: Zero-Shot (ZS), Few-Shot (FS), and Chain-of-Thought (CoT). As shown in Table 8, the performance variance across these paradigms remains below 1.6 points for all evaluated models, indicating that core capabilities are not significantly influenced by the prompting format.
Model $R_{\text{SOTA}}^{\text{model}}$ $S_{\text{model}}$ Eng. Econ. Edu. Law Lit. Mgmt. Sci. Hist. Med. Mil.
Open-source LLMs
DeepSeek-R1 97.40 91.23 9.47 9.43 9.27 9.37 8.83 9.37 9.03 9.53 8.50 8.43
DeepSeek-V3 96.47 90.36 9.30 9.57 8.93 9.23 8.60 9.13 8.97 9.47 8.83 8.33
Qwen-3-235B 96.43 90.32 9.23 9.43 9.03 9.50 8.23 9.43 8.97 9.17 8.73 8.60
Qwen-3-32B 92.22 86.38 8.43 9.10 8.57 9.10 7.77 9.47 8.67 9.30 7.70 8.27
QwQ-32B 94.49 88.54 8.30 9.46 9.23 9.33 7.83 9.46 8.65 9.27 8.57 8.43
GLM-4-32B 88.42 82.83 7.77 8.97 8.33 8.33 7.03 9.13 8.27 8.77 8.23 8.00
Qwen2.5-32B-Instruct 85.04 79.68 7.70 8.57 8.33 8.33 6.70 8.50 8.17 7.70 7.60 8.08
Qwen-Turbo-1101 83.68 78.41 7.97 8.37 8.03 8.23 6.40 8.50 8.10 7.50 7.27 8.05
Yi-34B-Chat 70.14 65.70 5.77 6.63 7.37 7.53 5.47 5.77 5.47 7.47 6.30 7.93
Nanbeige-plus 65.12 61.00 5.78 5.57 6.77 7.37 5.37 5.93 5.45 6.30 5.67 6.77
Baichuan2-13B-Chat 58.29 54.60 4.47 5.53 7.40 6.90 4.63 4.80 4.33 6.23 4.60 5.70
Llama-3.3-70B 60.84 57.00 5.80 6.90 5.63 5.70 5.47 5.70 6.30 4.70 4.87 5.93
Llama-3.2-90B-Vision 61.76 57.83 5.63 6.33 6.20 5.80 4.73 6.10 6.57 5.03 5.27 6.17
Qwen-plus 56.59 53.00 4.40 5.10 6.53 6.53 5.00 4.77 4.87 5.17 5.13 5.50
Qwen-turbo 55.76 52.23 4.10 6.07 6.63 6.43 4.43 4.53 4.97 5.27 4.37 5.43
Nanbeige-16B 55.44 51.93 4.37 5.30 6.50 6.30 3.97 4.70 4.07 5.90 4.73 6.10
Mixtral–8x7B-Instruct 51.67 48.40 4.27 5.47 6.47 6.40 3.13 4.50 5.07 3.57 4.37 5.17
ChatGLM2-6B 42.31 39.63 2.33 3.77 5.97 6.13 2.83 3.83 2.60 3.80 4.00 4.37
ziya_v1.1-13b 40.17 37.63 2.77 3.97 5.17 5.33 2.80 3.77 2.53 3.70 3.03 4.57
InternLM-Chat-7B 38.72 36.27 2.63 3.67 4.87 5.57 3.17 3.33 2.33 4.03 3.13 3.53
Phi-3-Medium-128K 36.94 34.60 2.27 4.17 3.70 4.23 2.87 4.50 3.57 3.20 2.27 3.83
Linly-Chinese-LLaMA-2-13B-hf 37.05 34.70 2.20 3.77 4.50 5.00 2.43 3.33 2.53 3.90 2.50 4.53
BELLE-Llama2-13B-chat-0.4M 36.26 33.97 2.57 3.07 4.93 4.73 2.83 3.80 2.43 3.33 2.40 3.87
Llama-2-7b-chat-hf 25.23 23.63 1.53 3.43 3.00 3.73 1.73 2.43 1.97 2.17 0.80 2.83
Closed-source LLMs
Doubao-1.5-Thinking-Pro 100.00 93.67 9.47 9.67 9.43 9.77 8.93 9.53 9.23 9.70 8.97 8.97
Gemini-2.5-Pro 97.22 91.07 9.20 9.47 9.20 9.30 8.43 9.63 9.07 9.40 8.50 8.87
Gemini-2.5-Pro-Thinking 97.15 91.00 9.13 9.50 9.37 9.47 8.40 9.63 9.20 9.27 8.30 8.73
Doubao-1.5-Pro 95.68 89.62 8.83 9.03 9.13 9.43 8.57 9.27 8.83 9.10 8.60 8.83
o1 93.36 87.45 8.90 9.30 8.67 8.77 7.73 9.27 8.90 8.97 8.17 8.77
Gemini-2.5-Flash-Thinking 92.71 86.87 8.67 9.27 8.70 9.00 7.80 8.93 8.90 9.00 8.03 8.57
Claude-Sonnet-4-Thinking 91.03 85.27 8.57 9.00 8.63 8.73 7.57 9.10 8.93 8.70 7.97 8.07
Claude-Sonnet-4 91.00 85.24 8.57 8.80 8.50 8.70 7.80 9.03 8.80 8.80 8.17 8.07
GPT-4o-search 89.40 83.74 8.27 8.77 8.43 8.67 7.77 8.80 8.20 8.73 8.27 7.83
GPT-4o 88.09 82.51 7.90 8.67 8.30 8.33 7.17 8.97 8.57 8.67 7.63 8.30
Gemini-1.5-Pro 85.91 80.48 8.13 8.45 8.30 8.37 7.04 8.17 8.43 8.50 7.48 7.60
o3-mini 84.13 78.80 7.97 8.60 8.30 8.20 6.73 8.57 8.53 7.17 7.03 7.70
Claude-3.5-Sonnet 83.36 78.10 7.97 8.53 8.27 7.93 7.03 8.50 8.00 7.57 6.70 7.60
o1-mini-2024-09-12 78.96 73.93 7.27 8.43 7.90 7.53 6.27 8.27 8.17 6.43 6.63 7.03
GPT4 Turbo 78.57 73.60 6.97 8.17 8.33 7.80 6.00 7.57 8.13 7.00 6.43 7.20
gpt-4-0125-preview 76.44 71.60 6.90 7.40 8.03 7.30 6.00 7.47 7.63 6.87 6.33 7.67
Baidu4.0 75.08 70.33 7.27 7.23 7.67 7.43 5.63 6.47 6.80 7.63 7.80 6.40
Baidu3.5 69.09 64.73 6.20 6.70 7.80 6.83 5.20 5.50 6.00 7.23 6.57 6.70
ChatGLM-pro 69.09 64.73 5.90 7.07 7.03 7.90 5.43 6.33 5.00 6.67 5.97 7.43
gpt4-0613 66.16 61.97 6.50 6.73 6.60 6.73 5.43 6.10 6.47 5.30 5.20 6.90
Spark-v3.0 65.62 61.47 5.77 6.50 7.27 7.30 5.70 5.90 5.03 6.50 5.23 6.27
Claude-3-Haiku 62.97 58.97 5.80 6.60 6.97 6.63 4.83 5.93 6.33 4.80 5.23 5.83
Gemini Pro 58.18 54.50 4.87 5.43 7.07 6.43 5.10 4.50 4.65 6.33 4.42 5.70
GPT-3.5-turbo 55.41 51.90 4.97 5.37 6.40 6.47 4.43 4.67 5.43 4.20 4.37 5.60
minimax-abab5 55.33 51.83 3.87 5.63 6.87 6.97 4.33 4.40 2.93 6.13 4.27 6.43
Table 7: Overall and Subject-Level Scores. $R_{\text{SOTA}}^{\text{model}}$ represents the relative score (0-100 scale) as defined in Equation (2), with Doubao-1.5-Thinking-Pro as the reference SOTA model. $S_{\text{model}}$ represents the absolute score (0-100 scale) as defined in Equation (1). Subject-level scores use a 10-point scale.
Model ZS FS CoT Mean Variance
Open-source LLMs
DeepSeek-R1 91.23 89.33 88.43 89.67 1.36
DeepSeek-V3 90.37 86.60 88.40 88.46 2.37
Qwen-3-235B 90.33 84.97 87.27 87.52 4.83
Qwen-3-32B 86.37 80.27 85.37 84.00 7.14
Closed-source LLMs
Doubao-1.5-Thinking-Pro 93.67 90.63 91.73 92.01 1.57
Gemini-2.5-Pro 91.07 87.57 90.60 89.74 2.41
Gemini-2.5-Pro-Thinking 91.00 87.87 90.23 89.70 1.78
Doubao-1.5-Pro 89.63 86.43 87.30 87.79 1.83
Claude-Sonnet-4-Thinking 85.27 85.60 86.87 85.91 0.48
Claude-Sonnet-4 85.23 84.77 86.23 85.41 0.37
GPT-4o 82.50 80.13 83.30 81.98 1.81
o1 87.43 86.07 86.87 86.79 0.31
o3-mini 78.80 76.97 79.67 78.48 1.27
Table 8: Model Capability under three prompting paradigms. ZS: Zero-Shot; FS: Few-Shot; CoT: Chain of Thought.
Appendix D Details in Experiment
This section provides additional experimental details demonstrating the robustness and stability of our ranking system. We present validation experiments across different sample sizes and show the consistency of our relative scoring methodology.
D.1 Sampling Validation
To verify the stability of our ranking system, we conducted evaluations across multiple sample sizes of n=1000 (in three rounds), 2000, and 4000 questions. The results in Table 9 demonstrate remarkable consistency in ranking order across all sample sizes.
Our relative scoring methodology produces smaller variance compared with absolute scoring approaches. The variance analysis reveals that top tier models show exceptional stability. DeepSeek-V3 exhibits a variance of 0.51 and O3-mini exhibits a variance of 0.95 while GPT-4o exhibits the highest variance of 1.63. Even the maximum variance represents less than two percent fluctuation indicating robust measurement precision.
The three independent one-thousand-sample runs demonstrate high reproducibility with models maintaining consistent relative positions across all test conditions. These findings validate that our ranking methodology captures stable model capabilities rather than random fluctuations.
D.2 LLM-as-Judge Validation
We calculated Cohen’s $\kappa$ coefficients between human evaluations and three LLM judges Doubao Gemini and GPT-4o across two evaluation rounds for thirteen models. As shown in Table 10 GPT-4o demonstrates superior performance with $\kappa$ values consistently above 0.90 with an average of 0.901 in Round 1 and 0.892 in Round 2 indicating almost perfect agreement.
The data reveal notable stability differences among judges. GPT-4o maintains consistently high agreement across both rounds with minimal variation while Doubao and Gemini exhibit more fluctuation between rounds. Specifically Doubao’s performance ranges from 0.232 to 0.745 across different models and Gemini exhibits even greater instability with some models showing dramatic drops between rounds—for example DeepSeek-V3 declines from 0.627 to 0.147.
In contrast Doubao and Gemini show lower overall agreement with average $\kappa$ values of 0.493 and 0.446 for Round 1 and Round 2 for Doubao and 0.494 and 0.400 for Gemini. Based on GPT-4o’s consistently high correlation with human evaluations and superior stability we selected it as our primary judge for reliable assessment.
Appendix E Implementation Details
This section provides detailed information about the annotation processes and evaluation procedures underlying our LLMEval-3 platform. We describe the expert involvement in data curation, validation processes, and the associated costs to ensure transparency and reproducibility.
E.1 Data Annotation Process
A total of 38 experts were engaged in data annotation and cleaning processes, with an average of more than 3 relevant specialists assigned to each discipline. For the annotation of original data, to mitigate fatigue-induced errors, annotation tasks for each expert were distributed across a 30–60 day period.
The cumulative remuneration disbursed to experts involved in data annotation and cleaning amounted to \$48,700. Ongoing investments are being allocated to further hire experts to expand the dataset.
E.2 Manual Evaluation Process
To validate our LLM-as-Judge approach, we conducted comprehensive human evaluation studies. A total of 18 experts participated in manual evaluation processes, with an average of about 2 relevant specialists assigned to each discipline. This expert-based validation ensures that our automated evaluation system maintains high agreement with human judgment standards.
The evaluation process involved multiple rounds of assessment across 13 representative models, with experts providing independent judgments that were subsequently compared against our LLM-based evaluation system using Cohen’s $\kappa$ coefficient.
E.3 Cost Analysis
The development and validation of LLMEval-3 required substantial investment in both human expertise and computational resources. We spent more than \$5,000 on using latest APIs of LLMs and deploying models for evaluation purposes. Additionally, \$10,000 was allocated for hiring qualified volunteers to conduct manual evaluations, ensuring rigorous validation of our automated assessment framework.
1000 Questions Larger Samples Statistics
Model Trial 1 Trial 2 Trial 3 2000 4000 Mean Variance
Doubao-1.5-Thinking-Pro 100.0 100.0 100.0 100.0 100.0 100.0 0.00
DeepSeek-V3 96.48 96.87 98.10 98.02 97.53 97.40 0.51
Qwen-3-32B 92.21 92.58 93.93 94.15 93.45 93.26 0.71
GPT-4o 88.08 90.21 91.50 90.69 90.48 90.19 1.63
o3-mini 84.13 85.31 86.69 85.56 86.18 85.57 0.95
Table 9: Ranking stability across different sample sizes. All scores are relative scores with Doubao-1.5-Thinking-Pro as reference (100.0). The three 1000-question trials demonstrate high reproducibility, with low variance indicating robust measurement precision.
Round 1 Round 2
Model Name Doubao Gemini GPT-4o Doubao Gemini GPT-4o
Open-source LLMs
DeepSeek-R1 0.527 0.662 0.960 0.451 0.251 0.977
DeepSeek-V3 0.326 0.627 0.949 0.366 0.147 0.800
Qwen-3-235B 0.486 0.468 0.818 0.232 0.496 0.931
Qwen-3-32B 0.560 0.390 0.886 0.414 0.271 0.872
Closed-source LLMs
Claude-Sonnet-4 0.309 0.186 0.826 0.402 0.677 0.909
Claude-Sonnet-4-Thinking 0.451 0.399 0.830 0.443 0.373 0.684
Doubao-1.5-Pro 0.629 0.388 0.950 0.383 0.451 0.915
Doubao-1.5-Thinking-Pro 0.707 0.786 0.993 0.745 0.324 0.920
Gemini-2.5-Pro 0.451 0.539 0.831 0.376 0.682 0.858
Gemini-2.5-Pro-Thinking 0.408 0.547 0.843 0.344 0.185 0.859
GPT-4o 0.447 0.507 0.925 0.553 0.417 0.962
o1 0.599 0.535 0.933 0.571 0.555 0.957
o3-mini 0.517 0.397 0.975 0.531 0.381 0.963
Mean 0.493 0.495 0.902 0.447 0.401 0.893
Table 10: Cohen’s $\kappa$ correlation coefficient between human evaluation and three large language model evaluations across two rounds.
Appendix F JWT Authentication Process
This section describes the detailed implementation of our JSON Web Token (JWT) authentication system, which forms the outer layer of our two-tier anti-cheating architecture. The JWT process ensures secure and authenticated access to our evaluation platform while preventing unauthorized access and session manipulation.
Our JWT implementation follows a standard three-phase protocol: token generation, transmission, and verification. Algorithm 1 outlines the complete JWT authentication workflow used in LLMEval-3.
Algorithm 1 JWT Authentication Process in LLMEval-3
1:Server-side Token Generation:
2: Generate a unique user identity (user_id)
3: Generate current timestamp and expiration time (exp)
4: Construct payload $\leftarrow$ {user_id, timestamp, exp, session_id, permissions}
5: Sign payload with server Secret using HMAC-SHA256 to generate JWT
6:return JWT to the authenticated user
7:
8:Client-side Request:
9: Include JWT in Authorization header: Bearer <token>
10: Send request to evaluation endpoint
11:
12:Server-side Verification:
13: Extract JWT from Authorization header
14: Verify JWT signature using server Secret
15: Parse payload and extract claims
16:if JWT signature is invalid then
17:  return HTTP 401 Unauthorized
18:end if
19:if current_time >> exp then
20:  return HTTP 401 Token Expired
21:end if
22:if user_id not found or permissions insufficient then
23:  return HTTP 403 Forbidden
24:end if
25:if session validation fails (e.g., concurrent sessions detected) then
26:  return HTTP 403 Session Invalid
27:end if
28: Allow evaluation operation to proceed
29: Log access attempt with user_id, timestamp, and session_id